What net force is needed to accelerate a 25 kg cartat 14 m/s2?
1
Expert's answer
2020-11-24T06:20:16-0500

Using the Newton's Second Law of Motion, we get:


"F_{net}=ma=25\ kg\cdot 14\ \dfrac{m}{s^2}=350\ N."

Answer:

"F_{net}=350\ N."